# Post-detection inference for sequential changepoint localization

### v2

@@ B.3.2 Known pre-change and composite post-change (Setting II)
B.3.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $200$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively. For change detection, we employ a weighted CUSUM-type detector (defined in (4.17)), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in Section B.3.2 is considered.
After detection, we construct confidence sets (as defined in (4.12)) for $T$ (assuming $T\neq 1$). We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs in Table B.4.
Table B.4: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ (4.6) and average detection delay (given no false detection) of the change detection method used (4.17) with threshold $A$, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
@@ B.3.3 Composite pre- and post-change (Setting III)
However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table B.2). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table B.2), exhibiting both higher coverage and larger size.
Appendix C Simulations for point estimates
In this section, we study the empirical performances of our point estimates of the changepoint defined in Equations 2.7, 3.5 and 4.6, which are the sequential analogs of the maximum likelihood estimators in the offline setting. Our empirical findings indicate that these estimators perform well in practice, even though we do not have theoretical guarantees on their properties.
C.1 Known pre- and post-change distributions
In this experiment, $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$ For detecting the change, we use the usual cumulative sum (CUSUM) and Shiryaev-Roberts (SR) procedures [23] as defined below:
$\tau_{\text{cusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak$ (C.1)
$\tau_{\text{sr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (C.2)
where $A$ is some fixed threshold. For our experiments, we set the value of $A$ at $10000$. Once the change is detected, we obtain $\hat{T}$, as defined in (2.7), based on observations up to the detection time. Fig. C.1 shows histograms of the values taken by $\hat{T}$ over $10000$ independent experiments.
Figure C.1: Histogram of $\hat{T}$, defined in (2.7), for known pre- and post-change over $10000$ independent experiments with $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$
C.2 Known pre-change and composite post-change distributions
In this experiment, $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta>0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
Here, we use the weighted CUSUM/SR detectors defined in (3.15) and (3.16). We considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. Fig. C.2 shows histograms of the values taken by $\hat{T}$ defined in (3.5) over $10000$ independent experiments.
Figure C.2: Histogram of $\hat{T}$, defined in (3.5), for known pre-change and unknown post-change over $10000$ independent experiments with $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta\neq 0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
C.3 Composite pre- and post-change distributions
In this experiment, $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively. We use the weighted CUSM/SR-type change detectors defined in (4.17) and (4.18).
In this experiment, we considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=0.6,d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ Fig. C.3 shows histograms of the values taken by $\hat{T}$ defined in (4.6) over $10000$ independent experiments. It is crucial to note that the RIPr-based estimator has a higher bias but lower variance.
Figure C.3: Histogram of $\hat{T}$, defined in (4.6), for unknown pre- and post-change over $10000$ independent experiments with $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
Appendix D A heuristic method for composite pre- or post-change
This section provides an alternative method for constructing confidence intervals for composite pre/post-change distributions. While this approach lacks theoretical guarantees, it performs slightly better in our experiments than our previous approach. Specifically, we propose a heuristic bootstrap-based approach that estimates the unknown distributions from the observed data and subsequently generates simulations from the estimated distributions.
D.1 Known pre-change and unknown post-change distributions
@@
For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined in (2.8)). Since the pre-change $F_{0}$ is known, we can generate $N$ many i.i.d. sequences under $F_{0}$ until a change is detected, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$. Now, the challenge in conducting a level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq t)$ test for each $t$ lies in the composite nature of $\mathcal{H}_{0,t}$, which prevents us from straightforwardly simulating data to determine a rejection threshold as we did in Algorithm 1. To address this, we employ a simple bootstrap-based approach. The main idea is to first obtain some point estimate of the changepoint (for example, a likelihood ratio based estimate, as defined in (3.5)). We denote the estimate as $\hat{T}$. Then we use the observations from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with known pre-change distribution $F_{0}$ and estimated post-change distribution $\hat{F}_{1}$ having changepoint at $t$.
@@
Input: $\alpha$, $N,B$, change detection algorithm $\mathcal{A}$ and data sequence $X_{1},X_{2},\cdots$ until change is detected at $\tau$ using $\mathcal{A}$
Output: A confidence set $\mathcal{C}$ for changepoint $T$
1 Compute a point estimator $\hat{T}$ for $T$, (eg., as defined in (3.5))
2 Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
3 for $i=1,\cdots,N$ do
4       Simulate $X_{1}^{i},X_{2}^{i}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first.
5
6 end for
@@
8 for $t=1,\cdots,\tau$ do
9       Compute some test statistics $M_{t}$ (eg., as defined in (3.7)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
10       for $j=1,\cdots,B$ do
11             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
12             Compute the same test statistics $M_{t}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}$
13
14       end for
15      if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
16            $\mathcal{C}=\mathcal{C}\cup\{t\}$
17       end if
18
19 end for
D.2 Composite pre- and post-change distributions
@@
For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined in (4.8)).
As in Algorithm 5, we employ a simple bootstrap-based approach, but here we need to estimate both pre- and post-change distributions. We first obtain some point estimate $\hat{T}$ of the changepoint $T$ (eg., the likelihhod-ratio-based estimate as defined in (4.6)). Then we use the observations from $1$ to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. For some $N\in\mathbb{N}$, we generate $N$ many i.i.d. sequences under $\hat{F}_{0}$ until a change is detected and denote the detection times as $\tau_{1},\cdots,\tau_{N}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with the estimated pre- and post-change distributions $\hat{F}_{0}$ and $\hat{F}_{1}$ respectively, having changepoint at $t$. For each sampled stream $j$, we calculate the test statistic $M_{t}^{j}$ to obtain $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if we have
@@
Input: $\alpha$, $N,B$, change detection algorithm $\mathcal{A}$ and data sequence $X_{1},X_{2},\cdots$ until change is detected at $\tau$ using $\mathcal{A}$
Output: A confidence set $\mathcal{C}$ for changepoint $T$
1 Compute some point estimator $\hat{T}$ for $T$ (eg., as defined in (4.6))
2 Obtain an estimate $\hat{F}_{0}$ of the (unknown) pre-change $F_{0}\in\mathcal{P}_{0}$ based on $X_{1},\cdots,X_{\hat{T}-1}$
3 Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
4 for $i=1,\cdots,N$ do
5       Simulate $X_{1}^{i},X_{2}^{i}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first.
6
7 end for
@@
9 for $t=1,\cdots,\tau$ do
10       Compute some test statistics $M_{t}$ (eg., as defined in (4.8)) based on $X_{1},\cdots,X_{\tau}$
11       for $j=1,\cdots,B$ do
12             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
13             Compute the same test statistics $M_{t}^{j}$ based on $X_{1}^{j},\cdots,X_{\tau_{j}}^{j}$
14
15       end for
16      if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
17            $\mathcal{C}=\mathcal{C}\cup\{t\}$
18       end if
19
20 end for
Algorithm 6 CI for unknown pre- and post-change
D.3 Experimental Results
D.3.1 Known pre-change and composite post-change (Setting II)
We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$. For detection, we employ a weighted CUSUM detector (3.15) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 5 to construct confidence sets. Table D.1 summarizes the results, including the average size and the conditional and unconditional coverage across $100$ independent runs. For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
Table D.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.15) with threshold $A=1000$.

### v3

@@ G.3.2 Known pre-change and composite post-change (Setting II)
G.3.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and 500500500, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively. For change detection, we employ a weighted CUSUM-type detector, where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, which is defined as
$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (G.7)
where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For the post-change parameter, the same discrete weight distribution as in Section G.3.2 is considered.
After detection, we construct confidence sets (as defined in Section C.2, assuming $T\neq 1$), with $\alpha=0.05,\beta=\gamma=0.025,B=N=100,L=\infty$. We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$, and delay (given no false detection) across 500500500 independent runs in Table G.9.
[…]
@@ G.3.3 Composite pre- and post-change (Setting III)
However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table G.7). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table G.7), exhibiting both higher coverage and larger size.
Appendix H A heuristic method for composite pre- or post-change
This section presents an alternative method for constructing confidence intervals for composite pre- and post-change distributions. Specifically, we propose a heuristic bootstrap-based approach that estimates the unknown distributions from the observed data and subsequently generates simulations from the estimated distributions.
H.1 Known pre-change and unknown post-change distributions
@@
For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$. Since the pre-change $F_{0}$ is known, we can generate $N$ many i.i.d. sequences under $F_{0}$ until a change is detected, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$. Now, the challenge in conducting a level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq t)$ test for each $t$ lies in the composite nature of $\mathcal{H}_{0,t}$, which prevents us from straightforwardly simulating data to determine a rejection threshold as we did in Algorithm 1. To address this, we employ a simple bootstrap-based approach. The main idea is to first obtain some point estimate of the changepoint (for example, a likelihood ratio based estimate, as defined earlier). We denote the estimate as $\hat{T}$. Then we use the observations from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with known pre-change distribution $F_{0}$ and estimated post-change distribution $\hat{F}_{1}$ having changepoint at $t$.
@@
2:A confidence set $\mathcal{C}$ for changepoint $T$
3:Compute a point estimator $\hat{T}$ for $T$
4:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
5:for $i=1$ to $N$ do
6:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
7:end for
[…]
9:for $t=1$ to $\tau$ do
10:  Compute some test statistics $M_{t}$ based on $\tau$ and $X_{1},\cdots,X_{\tau}$
11:  for $j=1$ to $B$ do
12:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
13:   Compute the same test statistics $M_{t}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}$
14:  end for
15:  if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
16:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
17:  end if
18:end for
H.2 Composite pre- and post-change distributions
@@
For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined earier).
As in Algorithm 4, we employ a simple bootstrap-based approach, but here we need to estimate both pre- and post-change distributions. We first obtain some point estimate $\hat{T}$ of the changepoint $T$. Then we use the observations from 111 to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. For some $N\in\mathbb{N}$, we generate $N$ many i.i.d. sequences under $\hat{F}_{0}$ until a change is detected and denote the detection times as $\tau_{1},\cdots,\tau_{N}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with the estimated pre- and post-change distributions $\hat{F}_{0}$ and $\hat{F}_{1}$ respectively, having changepoint at $t$. For each sampled stream $j$, we calculate the test statistic $M_{t}^{j}$ to obtain $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if we have
@@
Algorithm 5 CI for unknown pre- and post-change
@@
2:A confidence set $\mathcal{C}$ for changepoint $T$
3:Compute some point estimator $\hat{T}$ for $T$
4:Obtain an estimate $\hat{F}_{0}$ of the (unknown) pre-change $F_{0}\in\mathcal{P}_{0}$ based on $X_{1},\cdots,X_{\hat{T}-1}$
5:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
6:for $i=1$ to $N$ do
7:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
8:end for
[…]
10:for $t=1$ to $\tau$ do
11:  Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$
12:  for $j=1$ to $B$ do
13:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
14:   Compute the same test statistics $M_{t}^{j}$ based on $X_{1}^{j},\cdots,X_{\tau_{j}}^{j}$
15:  end for
16:  if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
17:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
18:  end if
19:end for
H.3 Experimental Results
H.3.1 Known pre-change and composite post-change (Setting II)
We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$. For detection, we employ a weighted CUSUM detector (B.11) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 4 to construct confidence sets. Table H.1 summarizes the results, including the average size and the conditional and unconditional coverage across 500500500 independent runs. For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
Table H.1: Setting II: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 4 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (B.11) with threshold $A=1000$.